OpenStudy (anonymous):

An investor wants to analyze the earnings of a mutual fund account. Four years ago, the value of the account was $16,000 and it is now worth $22,400. If the account is compared to a bank account paying interest that is compounded continuously, what interest rate rounded to the nearest hundredth of a percent would the bank account have to pay to match the mutual fund account’s earnings? (Assume the only deposit was to open the account. could you do this one with them plugged in the whole time?

OpenStudy (amistre64):

4 years ago, the value of the account was $16,000 and it is now worth $22,400. compounded continuously, find the rate of r A = Pe^(rt) ln(A/P) ------ = r t ln(16000/22400) --------------- = rate ; if your have questions about it, just ask 4

OpenStudy (anonymous):

-.0841?